### v1

# Using the rejection sampling for finding tests

Abstract
A new method based on the rejection sampling for finding statistical tests is proposed. This method is conceptually intuitive, easy to implement, and applicable for arbitrary dimension. To illustrate its potential applicability, three distinct empirical examples are presented: (1) examine the differences between group means of correlated (repeated) or independent samples, (2) examine if a mean vector equals to a specific fixed vector, and (3) investigate if samples come from a specific population distribution. The simulation examples indicate that the new test has similar statistical power as uniformly the most powerful (unbiased) tests. Moreover, these examples demonstrate that the new test is a powerful goodness-of-fit test. Keywords — Goodness-of-fit, joint hypothesis, likelihood ratio test, Monte Carlo $p$-value, rejection sampling, correlated samples
1 Introduction
Statistical hypothesis testing is a cornerstone of statistical inference. Numerous tests have been developed for various hypothesis testing problems, such as assessing statements about group means, testing for statistical independence (see, e.g., Reshef et al.,, 2018; Kuismin,, 2025), assess whether two samples came from the same underlying distribution (see, e.g., Giovanni and Franceschini,, 1987; Chen et al.,, 2025), test the normality of the data (see, e.g., Zhou and Shao,, 2014) or determining whether two categorical variables are associated, to mention just a few examples from the extensive body of statistical tests (see, e.g., Lehmann and Romano,, 2005; Kanji,, 2006).
The key goal in developing statistical tests is to create methods that can always distinguish genuine effects from those that arise purely by chance. In practice, however, a more realistic objective is to have a toolbox of tests that minimize the risk of erroneous conclusions about the population while maintaining a high probability of detecting true effects, provided that the specific characteristics of the problem at hand are carefully taken into account. The development of such statistical tests remains a fundamental problem in statistical research and continues to be an active area of investigation.
An addition to this toolbox is introduced in this paper. In particular, a new method for constructing statistical tests is proposed which is based on the rejection sampling commonly used to generate random samples from a probability distribution of interest. In sum, the current study aims to demonstrate how to use the probability of acceptance of the rejection sampling as a test statistic and provide a conceptual overview of the proposed approach. It is demonstrated that this method is versatile, easy to use, and exhibits high statistical power. The new method is used to derive tests for the following problems: (1) compare means of two independent or correlated variables; (2) examine the mean vector of a multivariate data of independent or correlated variables; and (3) investigate if samples come from a specific univariate or multivariate population distribution (goodness-of-fit). The performance of the tests derived using the new method is evaluated through Monte Carlo (MC) simulations to assess their statistical power. These examples suggest that this new framework demonstrates high statistical power as state-of-the-art tests, and in case (3) described above it exhibits even higher statistical power compared to the state-of-the-art tests.
2 Materials and method
Let $X=(x_{1},\ldots,x_{p})^{\top}$ be a $p$-dimensional real random vector and let $\textbf{X}=(X_{1},\ldots,X_{n})$ denote the data matrix. Following the notation of Casella and Berger, (2024), the general framework of statistical testing can be described as follows. Let $\theta$ denote a population parameter (vector) belonging to a parameter space $\Theta$. Let $\Theta_{0}$ be some subset of $\Theta$ and $\Theta_{0}^{c}$ is the complement of $\Theta_{0}$. In this paper $H_{0}$ and $H_{A}$ are used to denote the null and the alternative hypothesis, respectively. Consider testing $H_{0}:\theta\in\Theta_{0}$ against $H_{A}:\theta\in\Theta_{0}^{c}$.
[…]
A statistical hypothesis test is defined using a test statistic $T(X_{1},\ldots,X_{n})=T(\textbf{X})$. In most cases, the test says that $H_{0}$ should be rejected if $T(\textbf{X})$ is greater than (less than) or equal a certain constant $c$. This constant $c$ is called the rejection threshold, and the rejection region can be written as: $\{\textbf{X}\mid T(\textbf{X})\geq c\}$ where $c\in\mathbb{R}$.
There are different ways for finding test to assess if the observed data provides any evidence against $H_{0}$. Three of the most popular, well-known and studied methods are Wald test, the score test (also known as Lagrange multiplier), and the Likelihood Ratio (LR) test. From these methods the new framework presented in this paper is closely related to the LR test defined as follows:
Definition 1.
Let $\textbf{X}=(X_{1},\ldots,X_{n})$ be a random sample (data matrix) from a population with probability density function or probability mass function $f(\textbf{X}\mid\theta)$. The likelihood function is the joint density (or mass) function $L(\theta\mid\textbf{X})=f(\textbf{X}\mid\theta)$. The Likelihood Ratio test statistic for testing $H_{0}:\theta\in\Theta_{0}$ against $H_{A}:\theta\in\Theta_{0}^{c}$ is
$\lambda(\textbf{X})=\frac{\sup_{\Theta_{0}}L(\theta\mid\textbf{X})}{\sup_{\Theta}L(\theta\mid\textbf{X})}.$
A LR test is any test that has a rejection region of the form $\{\textbf{X}\mid\lambda(\textbf{X})\leq c\}$, where $0\leq c\leq 1$.
In the next section it is shown how the rejection sampling, also known as the accept-reject algorithm can be used to find statistical test. First, the basic principles of the AR algorithm are briefly reviewed to provide the necessary background, allowing the reader to follow the ideas presented in this paper.
2.1 The accept-reject algorithm
Rejection sampling, also known as the accept-reject algorithm (hereafter AR algorithm), is one of the conceptually simplest algorithms used to generate samples from arbitrary probability distributions of any dimension (see, e.g., Robert and Casella,, 2010, pp. 51–57). Let $f$ be a function presenting the functional form of the density of interest. For simplicity, $f$ is referred as the probability density function (pdf) of the distribution of interest; the algorithm is identical in case of probability mass function (pmf). $f$ is sometimes called as the target distribution. Use $X\sim f$ to indicate that the variable $X$ follows a specific but unknown multivariate probability distribution with density $f$. Let $g$ denote a known univariate or multivariate probability distribution with known pdf ($g$ is sometimes called as the proposal distribution). Let $D>0$ be a constant such that $f(X)\leq Dg(X)$ The AR algorithm can be outlined into the following four steps:
  1. 1.
    Generate $X_{i}=(x_{1i},\ldots,x_{pi})^{\top}\sim g$ and $u\sim Unif(0,1)$.
  2. 2.
    Compute $a=f(X_{i})/[Dg(X_{i})]$.
  3. 3.
    If $a>u$, accept $X_{i}$ as a sample from $f$.
  4. 4.
    Repeat steps 1.- 3. until the desired number of pseudo-random samples has been obtained.
The empirical efficiency of the AR algorithm can be evaluated using the so-called probability of acceptance, which is defined as follows: Let $\text{I}(x)$ denote the indicator function of the event $x>U$, where $U\sim\text{Unif}(0,1)$. It is assumed that $N\geq 1$ rounds of rejection sampling are performed. The probability of acceptance is given as the ratio of accepted samples to the total number of simulation rounds $N$:
$\rho=\frac{1}{N}\sum_{i=1}^{N}\text{I}\Bigl{(}\frac{f(X_{i})}{Dg(X_{i})}\Bigr{)}.$ (1)
The probability of acceptance, $0\leq\rho\leq 1$, can be used to evaluate the quality of the chosen proposal distribution $g$. It will be shown, that with small modifications the rejection sampling provides a very intuitive framework for deriving statistical tests.
To elaborate how the AR algorithm can be used to find tests, consider a a goodness-of-fit problem: $H_{0}:f(X_{i})=f_{0}(X_{i})$ for all $X_{i}$, $i=1,\ldots,n$ where $f$ is a population pdf and $f_{0}$ is a hypothesized pdf, e.g., multivariate normal density. Compare this to (1) and draw a parallel between $f_{0}$ and the target distribution of the AR algorithm. Under the null hypothesis, the density $g$ from which the data are drawn should agree with the density $f_{0}$. Instead of generating pseudo-random numbers from $g$, the observed samples $X_{1},\ldots,X_{n}$ are used as input to the algorithm. The likelihood ratios $f_{0}(X_{i})/g(X_{i})$, $i=1,\ldots,n$, can then be used to assess the degree of agreement between the hypothesized pdf $f_{0}$ and the data-dependent density $g$. In particular, the resulting probability of acceptance (1) should equal one if the observations are consistent with the null hypothesis ($f_{0}$). Intuitively, this corresponds the case that the target distribution $f_{0}$ and proposal distribution $g$ are equal; setting $D=1$ represents an uncommon case when using the AR algorithm but a logical choice when investigating if the data provides sufficient evidence against the specified theoretical distribution $f_{0}$.
However, $\rho$ depends on an external random variable $U$ meaning that the value of $\rho$ changes although the likelihood ratios $f(X_{i})/g(X_{i})$, $i=1,\ldots,n$ do not change. Therefore, instead of using (1) directly, the expected value of $\rho$ is used as the test statistic. The AR test statistic is defined as follows: Let $\text{I}(x)$ be the indicator function of the event $x>U$ where $U\sim Unif(0,1)$. Let $E(Y)$ be the expected value of a random variable $Y$. Let $T(\mathbf{X})=n^{-1}\sum_{i=1}^{n}\text{I}[f(X_{i})/g(X_{i})]$. The AR test statistic is the expected value of $T(\mathbf{X})$,
$\rho(\textbf{X})=E\left[T(\textbf{X})\right].$ (2)
The AR test is any test that has the rejection region $\{\textbf{X}\mid\rho(\textbf{X})\leq c\}$, where $0\leq c\leq 1$. From this point onward, notation is slightly abused: $T(\mathbf{X})$ is used to denote averages of indicator variables, each depending on differently defined likelihood ratios.
Here, two tests derived from the AR framework are considered. The first is a goodness-of-fit test, which can be defined as follows:
Definition 2.
Let $f_{0}$ be the claimed population pdf $f$ under $H_{0}:f=f_{0}$. Let $T(\mathbf{X})=n^{-1}\sum_{i=1}^{n}\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]$ be the mean of indicator variables depending on the ratios between the theoretical density $f_{0}$ and empirical density $\widehat{f}$, evaluated at the sample values. Use the AR test statistic (2) with rejection region $\{\textbf{X}\mid\rho(\textbf{X})\leq c\}$, where $0\leq c\leq 1$ for testing $H_{0}:f=f_{0}$ against $H_{A}:f\neq f_{0}$.
The second application of the AR test considered in this paper is to test values of a population parameter vector $\theta=(\theta_{1},\ldots,\theta_{q})^{\top}$, $q\geq 1$ defined as follows:
Definition 3.
[…]
The AR test is not limited to these tests; with small modifications it can be used in many testing problems. For example, in Kuismin, (2025) the author used a similar test to examine if two univariate random variables are statistically independent. The definition of the AR test is flexible: in principle, it can be defined with respect to different assumptions about the proposal distribution. For instance, while testing claims about the population covariance matrix (see, e.g., Bodnar et al.,, 2025), one could use the multivariate normal distribution, inverse Wishart distribution, the Lewandowski-Kurowicka-Joe distribution (Lewandowski et al.,, 2009), or even estimate the population distribution from data for the proposal and target distribution.
2.2 The distribution of the test statistic
Computing the distribution of $T(\textbf{X})$ by simulation is straightforward: By running the AR algorithm multiple times, the resulting values of $T(\textbf{X})$ can be used to approximate its distribution with respect to the data X. This allows evaluating the evidence provided by the AR test without resampling the data. The resulting interval estimates are credible intervals with a clear probabilistic interpretation. Moreover, it can be shown that the statistic $nT(\textbf{X})$ follows a Poisson binomial distribution when the goodness-of-fit test, as defined in Definition 2, is applied.
Let $b_{i}=\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]$, $i=1,\ldots,n$ as defined in Definition 2. From the properties of the Bernoulli distribution and the definition of the Poisson binomial distribution, it follows that
$nT(\textbf{X})=\sum_{i=1}^{n}b_{i}\sim\text{PoiBin}(\rho_{1},\ldots,\rho_{n}),$ (3)
where $\text{PoiBin}(\rho_{1},\ldots,\rho_{n})$ denotes a Poisson binomial distribution with parameters $\rho_{i}=P(b_{i}>U)=\text{min}\{1,b_{i}\}$, $i=1,\ldots,n$. When $\rho_{1}=\cdots=\rho_{n}=\rho^{\prime}$, the Poisson binomial distribution reduces to the binomial distribution $\text{Bin}(n,\rho^{\prime})$.
In Figure 1, the simulated and exact distributions of $nT(\textbf{X})$ are compared when the AR test is applied to test univariate normality (a detailed description of this example is provided in Subsection 2.5). In Figure 1, the Poisson binomial distribution and the simulated distribution of $nT(\textbf{X})$ are shown to be virtually identical. Consequently, the percentiles of the Poisson binomial distribution can be used to compute credible intervals for the statistic $T(\textbf{X})$ with respect to the data.
Figure 1: Probability mass function estimate of the statistic $nT(\textbf{X})$ based on 10000 simulation replicates (gray bars) and the Poisson binomial distribution (blue bars) when the AR test is used to test univariate normality (see Definition 2 and Section 2.5). Different panels represent different sample sizes: (A) 30; (B) 40; and (C) 70. The vertical dashed line illustrates the value $n\rho(\textbf{X})$.
The null distribution of test statistic $\rho(\textbf{X})$ must be specified in order to determine the rejection threshold $c$ and the $p$-value of the AR test. A natural choice is $c=1$, but this makes the test too liberal since $H_{0}$ would be rejected for any $\rho(\textbf{X})<1$. In addition, the null distribution is not well-defined in this case because the variance of the test statistic is zero. Thus, setting $c=1$ is not appropriate for all applications.
From the central limit theorem, it follows that the null distribution of $\rho(\textbf{X})$ is approximately normal. However, this normal approximation typically provides a crude estimate of the null distribution, since $\rho(\textbf{X})$ is close to one under $H_{0}$. For this reason, the null distribution is estimated using a simulation approach. In particular, a Monte Carlo procedure is used to determine the rejection threshold $c$ and, consequently, the $p$-value of the AR test (see, e.g., Howes,, 2023, for Monte Carlo significance testing). The following Monte Carlo significance testing procedure is used in this paper:
  1. 1.
    Generate $n$ independent and identically distributed (i.i.d.) values $\tilde{\textbf{X}}=(\tilde{X}_{1},\ldots,\tilde{X}_{n})$ from the target distribution $f_{0}$ under $H_{0}$.
  2. 2.
[…]
  3. 3.
    Repeat steps 1.- 2. $M$ times to get $\rho(\tilde{\textbf{X}})_{1},\ldots,\rho(\tilde{\textbf{X}})_{M}$ of the AR test statistic.
  4. 4.
    Use $\rho(\tilde{\textbf{X}})_{1},\ldots,\rho(\tilde{\textbf{X}})_{M}$ to estimate the null distribution of $\rho(\textbf{X})$ and calculate the Monte Carlo $p$-value, $p_{ar}=\sum_{m=1}^{M}[I(\rho(\tilde{\textbf{X}})_{m}\leq\rho(\textbf{X}))+1]/(M+1)$.
The percentile (rejection threshold) of the null distribution of $\rho(\mathbf{X})$ can be approximated by estimating the parameters of the Poisson binomial distribution with only a small number of Monte Carlo simulations. Table 1 shows percentiles of $\rho(\mathbf{X})$ based on 10000 Monte Carlo simulations and on the Poisson binomial distribution $\text{PoiBin}(\widehat{\rho}_{1},\ldots,\widehat{\rho}_{n})$, where $\widehat{\rho}_{i}$, $i=1,\ldots,n$, as defined in (3), are estimated under $H_{0}$ as the sample means from just 50 Monte Carlo replicates. The differences between percentiles obtained from the two approximations are negligible, and either method can be used without practical impact on inference as the sample size increases.
AR goodness-of-fit test AR group mean difference test
$n$ Percentile PoiBin approx. MC PoiBin approx. MC
10 0.01 0.600 0.586 0.600 0.517
0.05 0.700 0.719 0.700 0.660
0.10 0.800 0.777 0.700 0.719
30 0.01 0.800 0.803 0.767 0.763
0.05 0.867 0.852 0.800 0.820
0.10 0.867 0.878 0.833 0.849
60 0.01 0.883 0.863 0.850 0.841
0.05 0.900 0.899 0.867 0.876
0.10 0.917 0.915 0.883 0.893
Table 1: Percentiles of the empirical null distribution of $\rho(\textbf{X})$, estimated using the Poisson binomial distribution approximation and Monte Carlo (MC) simulations, for the AR-based goodness-of-fit test (Definition 2) and group mean difference test (Definition 3).
In the following sections, the empirical statistical power of three distinct AR tests is assessed through a comparative Monte Carlo simulation study, where the rejection threshold $c$ is determined by Monte Carlo simulations.
2.3 Comparing equality of group means
In this first example, the difference between group means is investigated, which represents one of the fundamental problems in statistical inference dating back to the pioneering work of Student, (1908). Let $\mu=(\mu_{1},\ldots,\mu_{p})^{\top}$ denote a population mean vector of length $p\geq 2$. Consider testing:
$H_{0}$ $:\mu_{1}=\mu_{2}=\ldots=\mu_{p}$
$H_{A}$ $:\mu_{i}\neq\mu_{j}\quad\text{ for some }i\text{ and }j\text{ where }i\neq j.$
Here the AR test defined in Definition 3 is applied for testing the hypothesis described above. In this example, it is demonstrated how incorporating certain distributional assumptions can improve the computational efficiency of the AR test. In particular, it is illustrated how the AR test can be constructed using a sufficient statistic. Here, the test is constructed based on the sample mean vector, which is a sufficient statistic for $\boldsymbol{\mu}$. The AR test for the sample mean vector is defined as follows: Let $\bar{X}=[\sum_{i=1}^{n}x_{j}^{i}/n]_{p\times 1}$ be the sample mean vector of length $p$. Denote $\mu_{0}=(\mu_{1}^{0},\ldots,\mu_{p}^{0})^{\top}$ be a parameter vector of length $p$ under $H_{0}$. Examine the value of the following statistic
$\textbf{t}=\sqrt{n}(\bar{X}-\mu_{0}).$ (4)
According to the multivariate central limit theorem, t converge in distribution to multivariate normal distribution $N(\textbf{0},\Sigma)$ under the null hypothesis where 0 is a zero vector of length $p$. Thus, the pdf of $N(\textbf{0},\Sigma)$ is used as the target distribution. Here a multivariate $t$-distribution is used as a proposal distribution to evaluate if the sample mean vector values differ from each other. Let $f_{n}\sim N(\textbf{0},\Sigma)$ and $g_{n}\sim t(n-1,\textbf{0},\Sigma)$ where $f_{n}$ and $g_{n}$ denote the pdf of the multivariate normal and multivariate $t$-distributions, respectively. Let $T(\textbf{t})=\text{I}[f_{n}(\textbf{t})/g_{n}(\textbf{t})]$ denote the indicator variable depending on the ratio between $f_{n}$ and $g_{n}$ under $H_{0}$. Clearly $f_{n}(\textbf{t})/g_{n}(\textbf{t})>0$ for all $\textbf{t}\in\mathbb{R}^{p}$ and $g_{n}(\textbf{t})\neq 0$. Under the null hypothesis set $\mu_{0}=[a]_{p\times 1}$ where $a=\sum_{j=1}^{p}\sum_{i=1}^{n}x_{i}^{j}/(pn)$. In this special case the AR test statistic simplifies to
$\rho(\textbf{X})=E\left[T(\textbf{t})>U\right]=P\left[T(\textbf{t})>U\right]=\min\{T(\textbf{t}),1\}.$ (5)
Now $f_{n}$ and $g_{n}$ depend on an unknown population parameter $\Sigma$. The sample covariance matrix $\widehat{\Sigma}=\sum_{i=1}^{n}(X_{i}-\bar{X})(X_{i}-\bar{X})^{\top}/(n-1)$ is employ in place of $\Sigma$.
The empirical power of the AR test is compared with the paired $t$-test, two-sample $t$-test, and the Likelihood Ratio (LR) test when $p=2$. According to the Neyman-Pearson Lemma, the LR test is uniformly the most powerful (UMP) test when considering a simple hypothesis $H_{0}:\mu_{d}=0$ against $H_{A}:\mu_{d}=\eta$ where $\mu_{d}=\mu_{1}-\mu_{2}$ and $\eta$ is the true difference between group means. The LR test is based on the assumption that all parameter values are known and does not involve parameter estimation. It is included here as a baseline test to illustrate the power that the optimal test could have when $p=2$ (see, e.g. Casella and Berger,, 2024, pp. 387–394). The sample size $n$ is set as 52 which is the minimal number of observations needed for the two sided paired $t$-test to have power at least $0.8$ when assuming that the group variances are equal and the effect size (Cohen’s D) is small (Cohen,, 1988).
For this simulation example, Let $X=(x_{1},x_{2})^{\top}\sim N(\mu,\Sigma)$ where $\mu=(\mu_{1},\mu_{2})^{\top}$ and $\Sigma=[\sigma_{ij}]$, $i,j\in\{1,2\}$ is a symmetric and positive definite matrix. In particular, consider $H_{0}:\mu_{1}=\mu_{2}$ against $H_{A}:\mu_{1}\neq\mu_{2}$. Here it is assumed that group variances are equal, $\sigma_{ii}=1$ and $\sigma_{ij}\in[-0.99,0.99]$. The power for each test is estimated when the difference between population mean parameters change. In particular, let $\eta=\mu_{1}-\mu_{2}$ when $\eta\in\{0.4,0.6,0.8\}$ at different values of the population correlation $\sigma_{ij}$. The empirical power of different tests is the ratio of significant tests of 10000 Monte Carlo samples at $\alpha=0.05$. Figure 2 illustrates the empirical power with respect to the population correlation.
Figure 2: Population correlation vs. the empirical power of the AR test (red solid line), Likelihood Ratio test (LR) (dotted purple line), paired $t$-test (solid red line), and independent sample $t$-test ($t$-test) (dashed blue line) at $\alpha=0.05$. Different panels illustrate the power vs. correlation when the true difference between two group means is: (A) 0.4; (B) 0.6; and (C) 0.8. Panel (D) illustrates the estimated Type I error associated with different tests. The predetermined significance level 0.05 is illustrated with horizontal dashed line in panel (D). Here $n=52$.
The power of the AR test is slightly lower compared to the paired $t$-test. However, from Figure 2 (D) one can see that the Type I error of the AR test as defined in (5) is below the predetermined level of significance. It appears that using a heavy-tailed distribution as the proposal distribution results in a conservative test. Figure 2 shows that when there is a strong positive, linear association between variables (groups) the statistical power of all methods increases.
The power estimates of the AR test are comparable to those of the two-sample $t$-test when the groups are independent. Figure 3 shows how the empirical power of independent sample tests change when the difference between group means change. The AR test shows practically equal statistical power compared to the $t$-test although the $t$-test used here assumes equal variance corresponding to the ground truth population parameter values.
Figure 3: Difference between two group means vs. the empirical power of the AR test (solid red line), Likelihood Ratio (LR) test (dashed green line), and Student’s $t$-test (dashed blue line) as a function of the true difference between group means when the sample size changes: (A) $n=26$; (B) $n=64$; and (C) $n=394$. The predetermined significance level 0.05 is illustrated with horizontal dashed line.
2.4 Testing if a population mean vector equals specific fixed vector
Let $\mu_{0}=(\mu_{0}^{1},\ldots,\mu_{0}^{p})^{\top}$ denote the assumption about the population mean vector $\mu$. Consider testing
$H_{0}$ $:\mu=\mu_{0}$
$H_{A}$ $:\mu_{j}\neq\mu_{j}^{0}\text{\, for some\,}j\in\{1,\ldots,p\}.$
[…]
In this example, $p=2$. Two powerful tests are selected for comparison: the LR test and the empirical LR test (EL) (see, e.g., Owen,, 2001). Their rejection regions are determined using the asymptotic distribution $\chi^{2}_{2}$. The EL test is computed with the R package mvhtests. The statistical power of the AR test, LR test, and EL is estimated as the ratio of significant tests among 10000 Monte Carlo samples. The empirical power of these tests is illustrated in Figure 4.
Figure 4: Population correlation vs. power function of the AR test using the population covariance (AR pop. cor) (red solid line), AR test using the sample covariance (AR sample cor) (green dashed line), empirical Likelihood Ratio test (EL) (dashed blue line), and Likelihood Ratio test (LR) (purple dashed line). Different panels illustrate the power vs. correlation when the true difference between two group means is: (A) 0.4; (B) 0.6; and (C) 0.8. Panel (D) illustrates the estimated Type I error associated with different tests. The predetermined significance level 0.05 is illustrated with horizontal dashed line in panel (D). Here $n=52$.
Figure 4 shows how the statistical power of the AR test, LR test, and the EL test are practically identical and all tests are very powerful. There is no difference whether the sample covariance or the population covariance matrix is used to compute the test statistic $\rho(\textbf{X})$.
2.5 A test for goodness-of-fit
Consider the investigation of whether a sample originates from a specific probability distribution. In this case, the hypothesis of the problem can be formulated as follows:
$H_{0}$ $:f(X_{i})=f_{0}(X_{i})\text{ for all }X_{i}\in\mathbb{R}^{p},$
$H_{A}$ $:f(X_{i})\neq f_{0}(X_{i})\text{ for some }X_{i}\in\mathbb{R}^{p},$
where $f$ is the theoretical population density distribution.
Here, the goodness-of-fit type AR test described in Definition 2 is employed to test the hypothesis described above. Let $\widehat{f}$ denote a density estimate constructed, for example, using the vine copula-based kernel density estimator of Nagler and Czado, (2016), which is considered an efficient method for kernel density estimation in high-dimensional data. Here, the R package kdevine is used to compute the density estimates $\widehat{f}$. The test statistic $T(\textbf{X})$ is defined as the averaged sum of indicator variables depending on ratio between $f_{0}(X_{i})$ and $\widehat{f}(X_{i})$, evaluated at the sample values,
$T(\textbf{X})=\frac{1}{n}\sum_{i=1}^{n}T(X_{i})=\frac{1}{n}\sum_{i=1}^{n}\text{I}\Bigl{[}f_{0}(X_{i})/\widehat{f}(X_{i})\Bigr{]}.$ (6)
The expected value of (6) can be compute applying the following Theorem (see also Kuismin,, 2025):
Theorem 1.
Let $r_{i}=f_{0}(X_{i})/\widehat{f}(X_{i})$. Let $S_{1}=\{r_{i}\mid r_{i}\geq 1\}$ and $S_{2}=\{r_{i}\mid r_{i}<1\}$ denote sets that contain values of ratios $r_{i}$, $i=1,\ldots,n$ which are greater than one or smaller than one, respectively. Let $S_{1}$ denote the cardinality of the set $S_{1}$. Let $T(\boldsymbol{X})$ be defined as in (6). Let $I(x)$ be the indicator function of the event $x>U$ where $U\sim Unif(0,1)$. The expected values of $T(\boldsymbol{X})$ is:
$E\left[T(\textbf{X})\right]=n^{-1}\left[\#S_{1}+\sum_{s_{2}\in S_{2}}s_{2}\right].$ (7)
Proof.
The expected value (7) follows from the properties of the indicator function, a uniformly distributed random variable, and the linearity of expectation. Let $U\sim\text{Unif}(0,1)$. Then $E[\text{I}(x)]=P(x>U)=1$ if $x\geq 1$, and $P(x>U)=x$ if $0<x<1$. From this it follows that
$E\left[T(\textbf{X})\right]$ $=E\Bigl{\{}n^{-1}\sum_{i=1}^{n}\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]\Bigr{\}}=n^{-1}\sum_{i=1}^{n}P\Bigl{[}f_{0}(X_{i})/\widehat{f}(X_{i})>U\Bigr{]}$
$=n^{-1}\left[\#S_{1}+\sum_{s_{2}\in S_{2}}s_{2}\right],$
which completes the proof. ∎
Thus, the AR test for the goodness-of-fit problem is given by
$\rho(\textbf{X})=n^{-1}\left[\#S_{1}+\sum_{s_{2}\in S_{2}}s_{2}\right].$ (8)
The AR test is compared with the Kolmogorov–Smirnov (KS) test (see, e.g., Giovanni and Franceschini,, 1987; Puritz et al.,, 2023), Cramér–von Mises (CVM) test (see, e.g., Chiu and Liu,, 2009), and Anderson-Darling test (AD) (Anderson and Darling,, 1952). While testing normality, a novel test proposed by Székely and Rizzo, (2005) (energy test (E)) is used. The AD and the energy tests are computed with R packages goftest and energy, respectively.
There are numerous distributions that can be tested using a goodness-of-fit test. Here it is investigated if a sample originate from a heavy-tailed univariate probability distribution with specific scale parameter value. In particular, it is examined wether the population distribution is a location-scale version of the $t$-distribution $t(df,\mu,\sigma)$ where $df$ is the number of degrees of freedom, $\mu$ is a location parameter, and $\sigma$ is a scale parameter. In this example, the null hypothesis is that the sample originates from $t(df=3,\mu=0,\sigma=\sigma_{0})$ distribution, where $\sigma_{0}=2.5$. Three different sample sizes are considered: $n=20,30$, and $50$. The empirical power of different tests is the ratio of significant tests of 10000 Monte Carlo samples at $\alpha=0.05$. The power of the AR, CVM, KS, and AD tests are illustrated in Figure 5.
Figure 5: Value of the scale parameter of the location-scale version of the t-distribution ($df=3$) vs. the empirical statistical power of the AR test (AR) (green dashed line), Cramér–von Mises test (CVM) (blue dashed line), the Kolmogorov–Smirnov test (KS) (purple dashed line), and Anderson-Darling test (AD) (solid red line). Different panels represent different sample sizes: (A) 20; (B) 30; and (C) 50. The vertical solid line corresponds to the value of the scale parameter under $H_{0}$, $\sigma_{0}=2.5$. The predetermined significance level $\alpha=0.05$ is illustrated with horizontal dashed line.
The results summarized in Figure 5 show that the power of the AR test is higher than that of the KS and CVM tests, even for the small sample sizes considered in this example. For $n=20$ and $n=30$, the power of the AD test is slightly higher than that of the AR test when $\sigma<\sigma_{0}$. However, as the sample size increases, the AR test appears to be the most powerful among those considered in this example.
The power of the AR test is also considered for assessing both univariate and multivariate normality. The null hypothesis is tested against four alternatives: the $t$-distribution $t(2)$, a normal mixture $0.5N(0,1)+0.5N(3,1)$, the logistic distribution, and the standard uniform distribution. The empirical power of each test is defined as the ratio of significant tests over 10000 Monte Carlo samples at $\alpha=0.05$. Estimates of the statistical power for the univariate normality test are summarized in Table 2 for sample sizes $20,30$ and $50$.
$n$ Alternative distribution AR KS CVM E AD
20 $N(0,1)$ 0.05 0.06 0.05 0.06 0.06
$t(2)$ 0.25 0.09 0.10 0.55 0.62
$0.5N(0,1)+0.5N(3,1)$ 0.95 0.96 0.96 0.15 1.00
logistic(0,1) 0.47 0.22 0.21 0.09 0.68
unif(0, 1) 1.00 1.00 1.00 0.15 1.00
30 $N(0,1)$ 0.04 0.05 0.05 0.05 0.05
$t(2)$ 0.37 0.11 0.11 0.67 0.73
$0.5N(0,1)+0.5N(3,1)$ 1.00 1.00 1.00 0.25 1.00
logistic(0,1) 0.72 0.34 0.34 0.13 0.87
unif(0,1) 1.00 1.00 1.00 0.28 1.00
50 $N(0,1)$ 0.05 0.05 0.05 0.06 0.05
$t(2)$ 0.59 0.15 0.16 0.85 0.85
$0.5N(0,1)+0.5N(3,1)$ 1.00 1.00 1.00 0.45 1.00
logistic(0,1) 0.96 0.53 0.58 0.14 0.97
unif(0,1 ) 1.00 1.00 1.00 0.53 1.00
Table 2: The empirical statistical power of the AR test (AR), the Kolmogorov–Smirnov test (KS), Cramér–von Mises test (CVM), energy test (E), and Anderson-Darling Test (AD) of univariate normality at $\alpha=0.05$. Here $N(0,1)$ corresponds to the null distribution.
As shown in Table 2, when the alternative distribution is the $t(2)$ distribution, the energy test and the AD test are more powerful than the AR test. Nevertheless, the AR test is still more powerful than KS and CVM tests. The AR test and the AD test achieve the highest statistical power when the alternative distribution is either a normal mixture or a logistic distribution. In the case of the logistic distribution, the empirical power of the AR test and the AD test is substantially higher than that of the KS, CVM, and energy tests.
When the AR test is applied as a test of multivariate normality, it is compared with five state-of-the-art multivariate procedures: the energy test, the test proposed by Zhou and Shao, (2014), referred to here as the $T_{n}$ test, test proposed by Henze and Zirkler (HZ test) (Henze and Zirkler,, 1990), Royston’s H test (Royston,, 1982), and test proposed by Villasenor-Alva and González-Estrada (VA-GE test) (Alva and Estrada,, 2009). The $T_{n}$ test, HZ test, Royston’s H test, and VA-GE test are computed with the R package mvnormalTest. The multivariate counterparts of the univariate distributions considered in the previous examples are used with dimension $p=3$. The empirical power of each test is defined as the ratio of significant tests over 1000 Monte Carlo samples at $\alpha=0.05$. Estimates of the power for the multivariate normality test are reported in Table 3.
$n$ Alternative distribution AR E $T_{n}$ HZ H VA-GE
20 $N(\textbf{0},I)$ 0.04 0.05 0.06 0.04 0.08 0.06
$t(2,\textbf{0},I)$ 0.05 0.84 0.82 0.78 0.83 0.64
$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 0.64 0.13 0.09 0.18 0.19 0.03
logistic(0, 1) 0.32 0.19 0.22 0.11 0.23 0.22
unif(0, 1)3 1.00 0.06 0.27 0.12 0.37 0.30
30 $N(\textbf{0},I)$ 0.06 0.04 0.05 0.05 0.08 0.05
$t(2,\textbf{0},I)$ 0.17 0.94 0.95 0.92 0.93 0.84
$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 0.93 0.25 0.16 0.40 0.35 0.04
logistic(0, 1) 0.81 0.24 0.30 0.16 0.26 0.29
unif(0, 1)3 1.00 0.12 0.54 0.29 0.73 0.67
50 $N(\textbf{0},I)$ 0.05 0.05 0.05 0.05 0.06 0.05
$t(2,\textbf{0},I)$ 0.59 1.00 1.00 0.99 0.99 0.98
$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 1.00 0.54 0.35 0.81 0.70 0.04
logistic(0, 1) 1.00 0.37 0.50 0.26 0.40 0.45
unif(0, 1)3 1.00 0.38 0.89 0.62 0.99 0.98
Table 3: The empirical statistical power of the AR test (AR), energy test (E), $T_{n}$ test, Henze and Zirkler test (HZ), Royston’s H test (H), and Villasenor-Alva and Gonzalez-Estrada test (VA-GE) of multivariate normality ($p=3$) at $\alpha=0.05$. Here $N(\textbf{0},I)$ corresponds to the null distribution. Here 0, 1, 3, and II denotes $(0,0,0)^{\top}$, $(1,1,1)^{\top}$, $(3,3,3)^{\top}$, and the identity matrix of size 33, respectively.
The results presented in Table 3 indicate that the power of the AR test is lower than that of the other tests when the alternative distribution is the multivariate $t$-distribution, although its power increases sharply (more than triples) as the sample size grows. In all other cases, the AR test is found to be the most powerful among the tests of multivariate normality considered.
3 Applications
[…]
3.1 Amyloid-beta
[…]
Group Comparison Diff. in Means $T(\textbf{X})$ 95% CI of $T(\textbf{X})$ $p$-value
NCI - mAD -425.03 0.589 [0.381; 0.810] 0.018
MCI - mAD -420.25 0.553 [0.333; 0.762] 0.009
NCI - MCI -4.78 1.00 [1.00; 1.00] 1.00
[…]
3.2 Reaction times
Reaction time (RT), typically measured in milliseconds as the time required to complete a task, is a widely used dependent variable in psychology (see, e.g., Whelan,, 2008; Boeck and Jeon,, 2019). RT distributions are almost always right-skewed, and thus asymmetric, heavy-tailed distributions are commonly employed to capture their shape. In this example, the AR goodness-of-fit test (8) is applied to RT data from Experiment 1 of Wagenmakers et al., (2008). In this experiment, participants were instructed to respond either as quickly or as accurately as possible. Here, the analysis focuses on the distribution of RTs from the “quick” responses.
Because a detailed analysis of the RTs does not contribute directly to this work, only the RTs from the participant with ID 1 are analyzed. Censored samples (uninterpretable responses, RT <180<180 ms, or RT >3>3 sec) and duplicated values were removed prior to analysis ($n=366$). A visual inspection of the data suggests that a shifted log-normal distribution provides a good fit to the reaction times. Figure 6 displays the histogram of the data along with the fitted shifted log-normal distribution, where the mean, standard deviation, and shift parameters were estimated using the R package EnvStats.
Figure 6: Histogram of the RTs for participant ID 1, along with the fitted shifted log-normal distribution (solid red line).
The value of the AR test statistic is $0.961$ under $H_{0}$ that the RTs follow a shifted log-normal distribution with the estimated parameter values. The corresponding Monte Carlo $p$-value is approximately $0.894$. The $95\%$ credible interval for the statistic $T(\textbf{X})$, computed from the percentiles of the Poisson binomial distribution (see (3)) is $[0.940;0.978]$. Taken together, these results suggest that the shifted log-normal distribution provides a close fit to the RTs considered in this example.
4 Conclusion
The AR algorithm possesses untapped potential for identifying powerful statistical tests. The AR tests illustrated in this paper achieve statistical power comparable to that of UMP tests commonly used in the literature. In particular, the goodness-of-fit test derived from the AR framework appears to outperform some state-of-the-art methods. The AR test is general and applicable to arbitrary dimensions, providing an easy-to-use test.
This paper only scratched the surface of the potentials of the method. With minor modifications the AR test can be utilized in many statistical inference problems, such as mixed data, categorical variables, the two-sample problem, the $K$-sample problem, among others. The introduction of these extensions is left for future research.
There are numerous factors that can affect the statistical power and Type I error of any test, including unequal or extremely small sample sizes, missing data, unequal group variances, skewness and kurtosis of underlying distributions, and the curse of dimensionality, to name just a few. Researchers encounter these challenges frequently in real-world data analysis. These important questions will be addressed in future studies.
Conflict of interest
The author declares no conflict of interest.
Data availability
The code to reproduce the examples, figures, and tables in this article are publicly available on GitHub at https://github.com/markkukuismin/Using-the-rejection-sampling-for-finding-tests-supplementary.
The Amyloid-$\beta$ data and the reaction time data are publicly available in the R packages Stat2Data and rtdists, respectively.

### v2

@@
 Statistical hypothesis testing is a cornerstone of statistical inference.
-Numerous tests have been developed for various hypothesis testing problems, such as assessing statements about group means, testing for statistical independence (see, e.g., Reshef et al.,, 2018; Kuismin,, 2025), assess whether two samples came from the same underlying distribution (see, e.g., Giovanni and Franceschini,, 1987; Chen et al.,, 2025), test the normality of the data (see, e.g., Zhou and Shao,, 2014) or determining whether two categorical variables are associated, to mention just a few examples from the extensive body of statistical tests (see, e.g., Lehmann and Romano,, 2005; Kanji,, 2006).
+Numerous tests have been developed for various hypothesis testing problems, such as assessing statements about group means, testing for statistical independence (see, e.g., Reshef et al., 2018; Kuismin, 2025), assess whether two samples came from the same underlying distribution (see, e.g., Giovanni and Franceschini, 1987; Chen et al., 2025), test the normality of the data (see, e.g., Zhou and Shao, 2014) or determining whether two categorical variables are associated, to mention just a few examples from the extensive body of statistical tests (see, e.g., Lehmann and Romano, 2005; Kanji, 2006).
 The key goal in developing statistical tests is to create methods that can always distinguish genuine effects from those that arise purely by chance.
@@
 These examples suggest that this new framework demonstrates high statistical power as state-of-the-art tests, and in case (3) described above it exhibits even higher statistical power compared to the state-of-the-art tests.
-2 Materials and method
-Let $X=(x_{1},\ldots,x_{p})^{\top}$ be a $p$-dimensional real random vector and let $\textbf{X}=(X_{1},\ldots,X_{n})$ denote the data matrix.
-Following the notation of Casella and Berger, (2024), the general framework of statistical testing can be described as follows.
-Let $\theta$ denote a population parameter (vector) belonging to a parameter space $\Theta$.
-Let $\Theta_{0}$ be some subset of $\Theta$ and $\Theta_{0}^{c}$ is the complement of $\Theta_{0}$.
-In this paper $H_{0}$ and $H_{A}$ are used to denote the null and the alternative hypothesis, respectively.
-Consider testing $H_{0}:\theta\in\Theta_{0}$ against $H_{A}:\theta\in\Theta_{0}^{c}$.
+2 Proposed framework
+Following Casella and Berger (2024), let $\theta$ denote a population quantity of interest (which may be a parameter vector or a more general functional) taking values in a space $\Theta$.
+Let $\Theta_{0}$ be some subset of $\Theta$, and let $\Theta_{0}^{c}$ denote its complement.
+Throughout this paper, $H_{0}$ and $H_{A}$ are used to denote the null and alternative hypotheses, respectively.
+In the parametric setting, the general format of null and alternative hypothesis is $H_{0}:\ \theta\in\Theta_{0}\quad\text{against}\quad H_{A}:\ \theta\in\Theta_{0}^{c}$.
+For example, one might be interested in testing $H_{0}:\ \theta=0\quad\text{against}\quad H_{A}:\ \theta\neq 0$ where $\theta$ is a population average parameter.
 […]
+Let $X=(x_{1},\ldots,x_{p})^{\top}$ be a $p$-dimensional real-valued random vector, and let $\mathbf{X}=(X_{1},\ldots,X_{n})$ denote the corresponding data matrix consisting of $n$ independent observations.
 A statistical hypothesis test is defined using a test statistic $T(X_{1},\ldots,X_{n})=T(\textbf{X})$.
@@
 2.1 The accept-reject algorithm
-Rejection sampling, also known as the accept-reject algorithm (hereafter AR algorithm), is one of the conceptually simplest algorithms used to generate samples from arbitrary probability distributions of any dimension (see, e.g., Robert and Casella,, 2010, pp.
+Rejection sampling, also known as the accept-reject algorithm (hereafter AR algorithm), is one of the conceptually simplest algorithms used to generate samples from arbitrary probability distributions of any dimension (see, e.g., Robert and Casella, 2010, pp.
 51–57).
@@
     Repeat steps 1.- 3. until the desired number of pseudo-random samples has been obtained.
-The empirical efficiency of the AR algorithm can be evaluated using the so-called probability of acceptance, which is defined as follows: Let $\text{I}(x)$ denote the indicator function of the event $x>U$, where $U\sim\text{Unif}(0,1)$.
+The empirical efficiency of the AR algorithm can be evaluated using the so-called probability of acceptance, which is defined as follows: Let $\text{I}(x)$ denote the indicator function/random variable of the event $x>U$, where $U\sim\text{Unif}(0,1)$.
 It is assumed that $N\geq 1$ rounds of rejection sampling are performed.
-The probability of acceptance is given as the ratio of accepted samples to the total number of simulation rounds $N$:
+The probability of acceptance is given as the ratio of accepted samples to the total number of simulation rounds $N$, and it can be estimated using the following estimator:
 $\rho=\frac{1}{N}\sum_{i=1}^{N}\text{I}\Bigl(\frac{f(X_{i})}{Dg(X_{i})}\Bigr).$ (1)
@@
 It will be shown, that with small modifications the rejection sampling provides a very intuitive framework for deriving statistical tests.
-To elaborate how the AR algorithm can be used to find tests, consider a a goodness-of-fit problem: $H_{0}:f(X_{i})=f_{0}(X_{i})$ for all $X_{i}$, $i=1,\ldots,n$ where $f$ is a population pdf and $f_{0}$ is a hypothesized pdf, e.g., multivariate normal density.
+2.2 Deriving a test using rejection sampling
+To elaborate how the AR algorithm can be used to find tests, consider a practical goodness-of-fit problem for multivariate distributions: $H_{0}:f(X_{i})=f_{0}(X_{i})$ for all $X_{i}$, $i=1,\ldots,n$ where $f$ is a population pdf and $f_{0}$ is a hypothesized pdf, e.g., multivariate normal density.
 Compare this to (1) and draw a parallel between $f_{0}$ and the target distribution of the AR algorithm.
@@
 Instead of generating pseudo-random numbers from $g$, the observed samples $X_{1},\ldots,X_{n}$ are used as input to the algorithm.
-The likelihood ratios $f_{0}(X_{i})/g(X_{i})$, $i=1,\ldots,n$, can then be used to assess the degree of agreement between the hypothesized pdf $f_{0}$ and the data-dependent density $g$.
-In particular, the resulting probability of acceptance (1) should equal one if the observations are consistent with the null hypothesis ($f_{0}$).
+The likelihood ratios $f_{0}(X_{i})/g(X_{i})$, $i=1,\ldots,n$, can then be used to assess the degree of agreement between the hypothesized pdf $f_{0}$ and density $g$.
+Let $g$ be a density estimate and assume, that $H_{0}$ is true.
+In this special case, the resulting probability of acceptance (1) should equal one if the observations are consistent with the null hypothesis ($f_{0}$).
 Intuitively, this corresponds the case that the target distribution $f_{0}$ and proposal distribution $g$ are equal; setting $D=1$ represents an uncommon case when using the AR algorithm but a logical choice when investigating if the data provides sufficient evidence against the specified theoretical distribution $f_{0}$.
@@
 Let $T(\mathbf{X})=n^{-1}\sum_{i=1}^{n}\text{I}[f(X_{i})/g(X_{i})]$.
-The AR test statistic is the expected value of $T(\mathbf{X})$,
-$\rho(\textbf{X})=E\left[T(\textbf{X})\right].$ (2)
+The AR test statistic is the expected value of $T(\mathbf{X})$ with respect to $U$,
+$\rho(\textbf{X})=E_{U}\left[T(\textbf{X})\right].$ (2)
 The AR test is any test that has the rejection region $\{\textbf{X}\mid\rho(\textbf{X})\leq c\}$, where $0\leq c\leq 1$.
 From this point onward, notation is slightly abused: $T(\mathbf{X})$ is used to denote averages of indicator variables, each depending on differently defined likelihood ratios.
-Here, two tests derived from the AR framework are considered. The first is a goodness-of-fit test, which can be defined as follows:
+In the next, two tests derived from the AR framework are considered.
+2.2.1 AR goodness-of-fit-test
+The first test is a goodness-of-fit test, which was already introduced in the previous section. More formally, it can be defined as follows:
 Definition 2.
 Let $f_{0}$ be the claimed population pdf $f$ under $H_{0}:f=f_{0}$.
-Let $T(\mathbf{X})=n^{-1}\sum_{i=1}^{n}\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]$ be the mean of indicator variables depending on the ratios between the theoretical density $f_{0}$ and empirical density $\widehat{f}$, evaluated at the sample values.
+Let $T(\mathbf{X})=n^{-1}\sum_{i=1}^{n}\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]$ be the mean of indicator variables depending on the ratios between the theoretical density $f_{0}$ and density estimate $\widehat{f}$, evaluated at the sample values.
 Use the AR test statistic (2) with rejection region $\{\textbf{X}\mid\rho(\textbf{X})\leq c\}$, where $0\leq c\leq 1$ for testing $H_{0}:f=f_{0}$ against $H_{A}:f\neq f_{0}$.
+2.2.2 AR test for a parameter vector
 The second application of the AR test considered in this paper is to test values of a population parameter vector $\theta=(\theta_{1},\ldots,\theta_{q})^{\top}$, $q\geq 1$ defined as follows:
@@
 […]
-The AR test is not limited to these tests; with small modifications it can be used in many testing problems.
-For example, in Kuismin, (2025) the author used a similar test to examine if two univariate random variables are statistically independent.
+The AR test is not limited to tests (2) and (3); with small modifications it can be used in many testing problems.
+For example, in Kuismin (2025) the author used a similar test to examine if two univariate random variables are statistically independent.
 The definition of the AR test is flexible: in principle, it can be defined with respect to different assumptions about the proposal distribution.
-For instance, while testing claims about the population covariance matrix (see, e.g., Bodnar et al.,, 2025), one could use the multivariate normal distribution, inverse Wishart distribution, the Lewandowski-Kurowicka-Joe distribution (Lewandowski et al.,, 2009), or even estimate the population distribution from data for the proposal and target distribution.
+For instance, while testing claims about the population covariance matrix (see, e.g., Bodnar et al., 2025), one could use the multivariate normal distribution, inverse Wishart distribution, the Lewandowski-Kurowicka-Joe distribution (Lewandowski et al., 2009), or even estimate the population distribution from data for the proposal and target distribution.
 2.3 The distribution of the test statistic
-Computing the distribution of $T(\textbf{X})$ by simulation is straightforward: By running the AR algorithm multiple times, the resulting values of $T(\textbf{X})$ can be used to approximate its distribution with respect to the data X.
+Estimating the distribution of $T(\textbf{X})$ by simulation is straightforward: By running the AR algorithm multiple times, the resulting values of $T(\textbf{X})$ can be used to approximate its distribution with respect to the data X.
 This allows evaluating the evidence provided by the AR test without resampling the data.
 The resulting interval estimates are credible intervals with a clear probabilistic interpretation.
-Moreover, it can be shown that the statistic $nT(\textbf{X})$ follows a Poisson binomial distribution when the goodness-of-fit test, as defined in Definition 2, is applied.
-Let $b_{i}=\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]$, $i=1,\ldots,n$ as defined in Definition 2.
-From the properties of the Bernoulli distribution and the definition of the Poisson binomial distribution, it follows that
+Moreover, it can be shown that the statistic $nT(\textbf{X})$ follows a Poisson binomial distribution.
+For simplicity, consider the case when the goodness-of-fit test, as defined in Definition 2, is applied.
+Let $b_{i}=\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]$, $i=1,\ldots,n$. From the properties of the Bernoulli distribution and the definition of the Poisson binomial distribution, it follows that
 $nT(\textbf{X})=\sum_{i=1}^{n}b_{i}\sim\text{PoiBin}(\rho_{1},\ldots,\rho_{n}),$ (3)
-where $\text{PoiBin}(\rho_{1},\ldots,\rho_{n})$ denotes a Poisson binomial distribution with parameters $\rho_{i}=P(b_{i}>U)=\text{min}\{1,b_{i}\}$, $i=1,\ldots,n$.
+where $\text{PoiBin}(\rho_{1},\ldots,\rho_{n})$ denotes a Poisson binomial distribution with parameters $\rho_{i}=P(b_{i}>U)=\text{min}\{f_{0}(X_{i})/\widehat{f}(X_{i}),1\}$, $i=1,\ldots,n$.
 When $\rho_{1}=\cdots=\rho_{n}=\rho^{\prime}$, the Poisson binomial distribution reduces to the binomial distribution $\text{Bin}(n,\rho^{\prime})$.
@@
 In addition, the null distribution is not well-defined in this case because the variance of the test statistic is zero.
-Thus, setting $c=1$ is not appropriate for all applications.
+Thus, setting $c=1$ is not appropriate for any application.
 From the central limit theorem, it follows that the null distribution of $\rho(\textbf{X})$ is approximately normal.
@@
 For this reason, the null distribution is estimated using a simulation approach.
-In particular, a Monte Carlo procedure is used to determine the rejection threshold $c$ and, consequently, the $p$-value of the AR test (see, e.g., Howes,, 2023, for Monte Carlo significance testing).
-The following Monte Carlo significance testing procedure is used in this paper:
+In particular, a Monte Carlo procedure is used to determine the rejection threshold $c$ and, consequently, the $p$-value of the AR test (see, e.g., Howes, 2023, for Monte Carlo significance testing).
+The following Monte Carlo significance testing procedure is used in this paper (where a small value of $\rho(\textbf{X})$ heuristically indicates evidence against $H_{0}$):
   1. 1.
@@
   3. 3.
-    Repeat steps 1.- 2. $M$ times to get $\rho(\tilde{\textbf{X}})_{1},\ldots,\rho(\tilde{\textbf{X}})_{M}$ of the AR test statistic.
+    Repeat steps 1 to 2 $M$ times to get $\rho(\tilde{\textbf{X}})_{1},\ldots,\rho(\tilde{\textbf{X}})_{M}$ of the AR test statistic.
   4. 4.
-    Use $\rho(\tilde{\textbf{X}})_{1},\ldots,\rho(\tilde{\textbf{X}})_{M}$ to estimate the null distribution of $\rho(\textbf{X})$ and calculate the Monte Carlo $p$-value, $p_{ar}=\sum_{m=1}^{M}[I(\rho(\tilde{\textbf{X}})_{m}\leq\rho(\textbf{X}))+1]/(M+1)$.
+    Use $\rho(\tilde{\textbf{X}})_{1},\ldots,\rho(\tilde{\textbf{X}})_{M}$ to estimate the null distribution of $\rho(\textbf{X})$ and calculate the Monte Carlo $p$-value, $p_{ar}=\sum_{m=1}^{M}[\text{I}(\rho(\tilde{\textbf{X}})_{m}\leq\rho(\textbf{X}))+1]/(M+1)$.
 The percentile (rejection threshold) of the null distribution of $\rho(\mathbf{X})$ can be approximated by estimating the parameters of the Poisson binomial distribution with only a small number of Monte Carlo simulations.
@@
 In the following sections, the empirical statistical power of three distinct AR tests is assessed through a comparative Monte Carlo simulation study, where the rejection threshold $c$ is determined by Monte Carlo simulations.
+It will be demonstrated numerically that, when $H_{0}$ is true, the probability that the AR tests incorrectly reject $H_{0}$ does not exceed the nominal significance level.
+Moreover, when a real effect is present, the power of the AR tests approaches one as the sample size increases.
+Each new AR test is also compared with state-of-the-art statistical tests.
 2.4 Comparing equality of group means
-In this first example, the difference between group means is investigated, which represents one of the fundamental problems in statistical inference dating back to the pioneering work of Student, (1908).
+This first example illustrates the parametric use of the AR test framework.
+In particular, the difference between group means is investigated, which represents one of the fundamental problems in statistical inference dating back to the pioneering work of Student (1908).
 Let $\mu=(\mu_{1},\ldots,\mu_{p})^{\top}$ denote a population mean vector of length $p\geq 2$.
@@
 In this example, it is demonstrated how incorporating certain distributional assumptions can improve the computational efficiency of the AR test.
-In particular, it is illustrated how the AR test can be constructed using a sufficient statistic.
+It is illustrated how the AR test can be constructed using a sufficient statistic.
 Here, the test is constructed based on the sample mean vector, which is a sufficient statistic for $\boldsymbol{\mu}$.
@@
 In this special case the AR test statistic simplifies to
-$\rho(\textbf{X})=E\left[T(\textbf{t})>U\right]=P\left[T(\textbf{t})>U\right]=\min\{T(\textbf{t}),1\}.$ (5)
+$\rho(\textbf{X})=E_{U}\left[T(\textbf{t})>U\right]=P\left[T(\textbf{t})>U\right]=\min\{T(\textbf{t}),1\}.$ (5)
 Now $f_{n}$ and $g_{n}$ depend on an unknown population parameter $\Sigma$.
@@
 It is included here as a baseline test to illustrate the power that the optimal test could have when $p=2$ (see, e.g.
-Casella and Berger,, 2024, pp.
+Casella and Berger, 2024, pp.
 387–394).
-The sample size $n$ is set as 52 which is the minimal number of observations needed for the two sided paired $t$-test to have power at least $0.8$ when assuming that the group variances are equal and the effect size (Cohen’s D) is small (Cohen,, 1988).
+The sample size $n$ is set as 52 which is the minimal number of observations needed for the two sided paired $t$-test to have power at least $0.8$ when assuming that the group variances are equal and the effect size (Cohen’s D) is small (Cohen, 1988).
 For this simulation example, Let $X=(x_{1},x_{2})^{\top}\sim N(\mu,\Sigma)$ where $\mu=(\mu_{1},\mu_{2})^{\top}$ and $\Sigma=[\sigma_{ij}]$, $i,j\in\{1,2\}$ is a symmetric and positive definite matrix.
-In particular, consider $H_{0}:\mu_{1}=\mu_{2}$ against $H_{A}:\mu_{1}\neq\mu_{2}$.
+Consider the hypothesis test $H_{0}:\mu_{1}=\mu_{2}$ against $H_{A}:\mu_{1}\neq\mu_{2}$.
 Here it is assumed that group variances are equal, $\sigma_{ii}=1$ and $\sigma_{ij}\in[-0.99,0.99]$.
 The power for each test is estimated when the difference between population mean parameters change.
-In particular, let $\eta=\mu_{1}-\mu_{2}$ when $\eta\in\{0.4,0.6,0.8\}$ at different values of the population correlation $\sigma_{ij}$.
+Let $\eta=\mu_{1}-\mu_{2}$ when $\eta\in\{0.4,0.6,0.8\}$ at different values of the population correlation $\sigma_{ij}$.
 The empirical power of different tests is the ratio of significant tests of 10000 Monte Carlo samples at $\alpha=0.05$.
@@
 In this example, $p=2$.
-Two powerful tests are selected for comparison: the LR test and the empirical LR test (EL) (see, e.g., Owen,, 2001).
+Two powerful tests are selected for comparison: the LR test and the empirical LR test (EL) (see, e.g., Owen, 2001).
 Their rejection regions are determined using the asymptotic distribution $\chi^{2}_{2}$.
@@
 Here, the goodness-of-fit type AR test described in Definition 2 is employed to test the hypothesis described above.
-Let $\widehat{f}$ denote a density estimate constructed, for example, using the vine copula-based kernel density estimator of Nagler and Czado, (2016), which is considered an efficient method for kernel density estimation in high-dimensional data.
-Here, the R package kdevine is used to compute the density estimates $\widehat{f}$.
+Let $\widehat{f}$ denote a density estimate constructed, for example, using kernel density estimation.
+Here, the R package Rfast2 using a high-degree memory and computationally optimized kernel functions is used to compute the density estimates $\widehat{f}$ Tsagris and Papadakis (2018).
 The test statistic $T(\textbf{X})$ is defined as the averaged sum of indicator variables depending on ratio between $f_{0}(X_{i})$ and $\widehat{f}(X_{i})$, evaluated at the sample values,
 $T(\textbf{X})=\frac{1}{n}\sum_{i=1}^{n}T(X_{i})=\frac{1}{n}\sum_{i=1}^{n}\text{I}\Bigl[f_{0}(X_{i})/\widehat{f}(X_{i})\Bigr].$ (6)
-The expected value of (6) can be compute applying the following Theorem (see also Kuismin,, 2025):
+The expected value of (6) can be compute applying the following Theorem (see also Kuismin, 2025):
 Theorem 1.
@@
 The expected values of $T(\boldsymbol{X})$ is:
-$E\left[T(\textbf{X})\right]=n^{-1}\left[\#S_{1}+\sum_{s_{2}\in S_{2}}s_{2}\right].$ (7)
+$E_{U}\left[T(\textbf{X})\right]=n^{-1}\left[\#S_{1}+\sum_{s_{2}\in S_{2}}s_{2}\right].$ (7)
 Proof.
@@
 From this it follows that
-$E\left[T(\textbf{X})\right]$ $=E\Bigl{\{}n^{-1}\sum_{i=1}^{n}\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]\Bigr{\}}=n^{-1}\sum_{i=1}^{n}P\Bigl{[}f_{0}(X_{i})/\widehat{f}(X_{i})>U\Bigr{]}$
+$E_{U}\left[T(\textbf{X})\right]$ $=E_{U}\Bigl\{n^{-1}\sum_{i=1}^{n}\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]\Bigr\}=n^{-1}\sum_{i=1}^{n}P\Bigl[f_{0}(X_{i})/\widehat{f}(X_{i})>U\Bigr]$
 $=n^{-1}\left[\#S_{1}+\sum_{s_{2}\in S_{2}}s_{2}\right],$
@@
 $\rho(\textbf{X})=n^{-1}\left[\#S_{1}+\sum_{s_{2}\in S_{2}}s_{2}\right].$ (8)
-The AR test is compared with the Kolmogorov–Smirnov (KS) test (see, e.g., Giovanni and Franceschini,, 1987; Puritz et al.,, 2023), Cramér–von Mises (CVM) test (see, e.g., Chiu and Liu,, 2009), and Anderson-Darling test (AD) (Anderson and Darling,, 1952).
-While testing normality, a novel test proposed by Székely and Rizzo, (2005) (energy test (E)) is used.
+The AR test is compared with the Kolmogorov–Smirnov (KS) test (see, e.g., Giovanni and Franceschini, 1987; Puritz et al., 2023), Cramér–von Mises (CVM) test (see, e.g., Chiu and Liu, 2009), and Anderson-Darling test (AD) (Anderson and Darling, 1952).
+While testing normality, a novel test proposed by Székely and Rizzo (2005) (energy test (E)) is used.
 The AD and the energy tests are computed with R packages goftest and energy, respectively.
@@
 The results summarized in Figure 5 show that the power of the AR test is higher than that of the KS and CVM tests, even for the small sample sizes considered in this example.
-For $n=20$ and $n=30$, the power of the AD test is slightly higher than that of the AR test when $\sigma<\sigma_{0}$.
-However, as the sample size increases, the AR test appears to be the most powerful among those considered in this example.
+The power of the AR test is slightly higher than that of the AD test when $\sigma<\sigma_{0}$.
+As the sample size increases, the AR test appears to be the most powerful among those considered in this example.
 The power of the AR test is also considered for assessing both univariate and multivariate normality.
@@
 $n$ Alternative distribution AR KS CVM E AD
-20 $N(0,1)$ 0.05 0.06 0.05 0.06 0.06
-$t(2)$ 0.25 0.09 0.10 0.55 0.62
-$0.5N(0,1)+0.5N(3,1)$ 0.95 0.96 0.96 0.15 1.00
-logistic(0,1) 0.47 0.22 0.21 0.09 0.68
+20 $N(0,1)$ 0.04 0.05 0.05 0.05 0.05
+$t(2)$ 0.27 0.09 0.10 0.51 0.60
+$0.5N(0,1)+0.5N(3,1)$ 0.99 0.97 0.96 0.15 1.00
+logistic(0,1) 0.65 0.23 0.23 0.10 0.72
 unif(0, 1) 1.00 1.00 1.00 0.15 1.00
-30 $N(0,1)$ 0.04 0.05 0.05 0.05 0.05
-$t(2)$ 0.37 0.11 0.11 0.67 0.73
-$0.5N(0,1)+0.5N(3,1)$ 1.00 1.00 1.00 0.25 1.00
-logistic(0,1) 0.72 0.34 0.34 0.13 0.87
+30 $N(0,1)$ 0.05 0.05 0.05 0.05 0.05
+$t(2)$ 0.36 0.10 0.11 0.67 0.72
+$0.5N(0,1)+0.5N(3,1)$ 1.00 1.00 1.00 0.24 1.00
+logistic(0,1) 0.81 0.33 0.34 0.12 0.85
 unif(0,1) 1.00 1.00 1.00 0.28 1.00
-50 $N(0,1)$ 0.05 0.05 0.05 0.06 0.05
-$t(2)$ 0.59 0.15 0.16 0.85 0.85
-$0.5N(0,1)+0.5N(3,1)$ 1.00 1.00 1.00 0.45 1.00
-logistic(0,1) 0.96 0.53 0.58 0.14 0.97
+50 $N(0,1)$ 0.05 0.05 0.05 0.05 0.05
+$t(2)$ 0.55 0.14 0.14 0.85 0.86
+$0.5N(0,1)+0.5N(3,1)$ 1.00 1.00 1.00 0.44 1.00
+logistic(0,1) 0.96 0.52 0.57 0.14 0.96
 unif(0,1 ) 1.00 1.00 1.00 0.53 1.00
@@
 In the case of the logistic distribution, the empirical power of the AR test and the AD test is substantially higher than that of the KS, CVM, and energy tests.
-When the AR test is applied as a test of multivariate normality, it is compared with five state-of-the-art multivariate procedures: the energy test, the test proposed by Zhou and Shao, (2014), referred to here as the $T_{n}$ test, test proposed by Henze and Zirkler (HZ test) (Henze and Zirkler,, 1990), Royston’s H test (Royston,, 1982), and test proposed by Villasenor-Alva and González-Estrada (VA-GE test) (Alva and Estrada,, 2009).
+When the AR test is applied as a test of multivariate normality, it is compared with five state-of-the-art multivariate procedures: the energy test, the test proposed by Zhou and Shao (2014), referred to here as the $T_{n}$ test, test proposed by Henze and Zirkler (HZ test) (Henze and Zirkler, 1990), Royston’s H test (Royston, 1982), and test proposed by Villasenor-Alva and González-Estrada (VA-GE test) (Alva and Estrada, 2009).
 The $T_{n}$ test, HZ test, Royston’s H test, and VA-GE test are computed with the R package mvnormalTest.
@@
 20 $N(\textbf{0},I)$ 0.04 0.05 0.06 0.04 0.08 0.06
-$t(2,\textbf{0},I)$ 0.05 0.84 0.82 0.78 0.83 0.64
-$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 0.64 0.13 0.09 0.18 0.19 0.03
-logistic(0, 1) 0.32 0.19 0.22 0.11 0.23 0.22
+$t(2,\textbf{0},I)$ 0.11 0.84 0.82 0.78 0.83 0.64
+$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 0.99 0.13 0.09 0.18 0.19 0.03
+logistic(0, 1) 0.76 0.19 0.22 0.11 0.23 0.22
 unif(0,1)3 1.00 0.06 0.27 0.12 0.37 0.30
 30 $N(\textbf{0},I)$ 0.06 0.04 0.05 0.05 0.08 0.05
-$t(2,\textbf{0},I)$ 0.17 0.94 0.95 0.92 0.93 0.84
-$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 0.93 0.25 0.16 0.40 0.35 0.04
-logistic(0, 1) 0.81 0.24 0.30 0.16 0.26 0.29
+$t(2,\textbf{0},I)$ 0.09 0.94 0.95 0.92 0.93 0.84
+$0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 1.00 0.25 0.16 0.40 0.35 0.04
+logistic(0, 1) 0.90 0.24 0.30 0.16 0.26 0.29
 unif(0,1)3 1.00 0.12 0.54 0.29 0.73 0.67
-50 $N(\textbf{0},I)$ 0.05 0.05 0.05 0.05 0.06 0.05
-$t(2,\textbf{0},I)$ 0.59 1.00 1.00 0.99 0.99 0.98
+50 $N(\textbf{0},I)$ 0.07 0.05 0.05 0.05 0.06 0.05
+$t(2,\textbf{0},I)$ 0.09 1.00 1.00 0.99 0.99 0.98
 $0.5N(\textbf{0},I)+0.5N(\textbf{3},I)$ 1.00 0.54 0.35 0.81 0.70 0.04
-logistic(0, 1) 1.00 0.37 0.50 0.26 0.40 0.45
+logistic(0, 1) 0.98 0.37 0.50 0.26 0.40 0.45
 unif(0,1)3 1.00 0.38 0.89 0.62 0.99 0.98
@@
 Here 0, 1, 3, and II denotes $(0,0,0)^{\top}$, $(1,1,1)^{\top}$, $(3,3,3)^{\top}$, and the identity matrix of size 33, respectively.
-The results presented in Table 3 indicate that the power of the AR test is lower than that of the other tests when the alternative distribution is the multivariate $t$-distribution, although its power increases sharply (more than triples) as the sample size grows.
+The results presented in Table 3 indicate that the power of the AR test is lower than that of the other tests when the alternative distribution is the multivariate $t$-distribution.
 In all other cases, the AR test is found to be the most powerful among the tests of multivariate normality considered.
+All of these examples show that the AR test maintains the Type I error at a pre-specified level of significance.
+The power of the AR test depends on the density estimation method.
+For example, instead of using the kernel density estimator from the R package Rfast2, one viable option is to employ the vine-copula-based kernel density estimator of Nagler and Czado (2016), which is considered an efficient method for kernel density estimation in high-dimensional settings.
+Using this estimator improves the power of the AR test when the alternative distribution is a heavy-tailed t-distribution.
+However, in the other cases considered here, the statistical power decreases slightly (results not shown).
 3 Applications
@@
 3.2 Reaction times
-Reaction time (RT), typically measured in milliseconds as the time required to complete a task, is a widely used dependent variable in psychology (see, e.g., Whelan,, 2008; Boeck and Jeon,, 2019).
+Reaction time (RT), typically measured in milliseconds as the time required to complete a task, is a widely used dependent variable in psychology (see, e.g., Whelan, 2008; Boeck and Jeon, 2019).
 RT distributions are almost always right-skewed, and thus asymmetric, heavy-tailed distributions are commonly employed to capture their shape.
-In this example, the AR goodness-of-fit test (8) is applied to RT data from Experiment 1 of Wagenmakers et al., (2008).
+In this example, the AR goodness-of-fit test (8) is applied to RT data from Experiment 1 of Wagenmakers et al.
+(2008).
 In this experiment, participants were instructed to respond either as quickly or as accurately as possible.
@@
 A visual inspection of the data suggests that a shifted log-normal distribution provides a good fit to the reaction times.
-Figure 6 displays the histogram of the data along with the fitted shifted log-normal distribution, where the mean, standard deviation, and shift parameters were estimated using the R package EnvStats.
-Figure 6: Histogram of the RTs for participant ID 1, along with the fitted shifted log-normal distribution (solid red line).
+Figure 6 displays the histogram of the data along with the fitted shifted log-normal distribution, where the mean, standard deviation, and shift parameters were estimated using the R package EnvStats, as well as the fitted normal distribution, for which the mean and standard deviation were estimated from the data.
+Figure 6: Histogram of the RTs for participant ID 1, along with the fitted shifted log-normal distribution (solid red line), and normal distribution (dashed blue line).
 The value of the AR test statistic is $0.961$ under $H_{0}$ that the RTs follow a shifted log-normal distribution with the estimated parameter values.
-The corresponding Monte Carlo $p$-value is approximately $0.894$.
-The $95\%$ credible interval for the statistic $T(\textbf{X})$, computed from the percentiles of the Poisson binomial distribution (see (3)) is $[0.940;0.978]$.
-Taken together, these results suggest that the shifted log-normal distribution provides a close fit to the RTs considered in this example.
+The AR test statistic is $0.758$ under $H_{0}$ that the RTs follow a normal distribution.
+The corresponding Monte Carlo $p$-value is approximately $0.894$ and $0.001$, respectively.
+The $95\%$ credible interval for the statistic $T(\textbf{X})$, computed from the percentiles of the Poisson binomial distribution (see (3)) is $[0.940;0.978]$ and $[0.713;0.801]$, respectively.
+Taken together, these results suggest that the shifted log-normal distribution provides a close fit to the RTs considered in this example, offering a substantially better approximation than the normality assumption.
+Note that, although using estimated parameter values in hypothesis testing is questionable, this example illustrates how the AR test can be applied to evaluate different distributional assumptions rather than relying solely on the visual inspection of the data..
 4 Conclusion
@@
 These important questions will be addressed in future studies.
-Conflict of interest
-The author declares no conflict of interest.
 Data availability